What was the average speed of the man in m/min? (68 m/min)
A boy runs at 5.0 m/s [30° S of W] for 2.5 minutes and then he turns and runs at
3.0 m/s [40° S of E] for 4.5 minutes.
A. What was his average speed?
B. What was his displacement?
1
Expert's answer
2019-10-03T09:29:17-0400

A. The average speed:


"v=\frac{s}{t}=\frac{v_1t_1+v_2t_2}{t_1+t_2}"

"v=\frac{(5)2.5+(3)4.5}{2.5+4.5}=3.7\frac{m}{s}"

B.


"\bold{d_1}=((5\cdot 2.5\cdot 60)\cos{30\degree},(5\cdot 2.5\cdot60)\sin{30\degree})=(649.5,375)m"

"\bold{d_2}=(-(3\cdot 4.5\cdot 60)\cos{40\degree},(3\cdot 4.5\cdot 60)\sin{40\degree})=(-620.5,520.7)m"

"\bold{d}=(649.5-620.5,375+520.7)m=(29,895.7)m"

"\bold{d}=896\ m \ 1.9\degree W of S"